# Sparse and Faithful Explanations Without Sparse Models

### 04 Mar 2024 v2

@@ 4.2 Individual-based SEV Optimization
The volume-based SEV loss above approximates queries as arising uniformly over $\mathcal{X}$, which might not be true. Making the less restrictive assumption that the distribution of a training dataset is close to that of test data, we can instead focus on minimizing the average training SEV. The difficulty here lies in the combinatorial number of steps required to calculate SEV. However, we have found that most trained classifiers yield low average SEVs (roughly $<3$) on real datasets. With the majority of training data having low SEV, we can design the optimization objective to penalize samples with ${}^{+}$ not equal to 1. To do this, we maximize the fraction of points where ${}^{+}$ is exactly 1. Given a class $\mathcal{F}$ of classification models that output the probability of belonging to the positive class, the loss for minimizing ${}^{+}$ on a training set is:
@@
$\frac{-1}{n^{+}}\sum_{i=1}^{n^{+}}\min\left(\max_{j=1,\dots,p}g(\boldsymbol{e}_{j}\odot\boldsymbol{x}_{i}+(\mathbf{1}-\boldsymbol{e}_{j})\odot\tilde{\boldsymbol{x}}),~{}\mathcal{T}\right)$
where $\boldsymbol{x}_{i}$ is the $i$-th query in the training data, $n^{+}$ represents the number of positively predicted queries in the dataset, and $\boldsymbol{e}_{j}$ denotes the vector with a 1 in the $j^{th}$ coordinate and 0’s elsewhere, and $g:\mathbb{R}^{p}\to[0,1]$ is the continuous estimator associated with classifier $f$ (e.g., in logistic regression, $g$ is the probability of the class of interest), and $\mathcal{T}$ is the threshold value for classification, which is usually 0.5 for balanced data. Here, we consider all vertices 1 step away from the reference and see whether at least one of them is predicted as positive (i.e., ${}^{+}$=1). The min operation taken with the threshold $\mathcal{T}$ is used to fix the value of the objective when the prediction is positive since we do not care where on the positive side of the decision boundary a prediction is, as long as it is positive. Similarly, we propose the following loss term for minimizing ${}^{-}$:
@@
We call these individual-based SEV losses ${}^{+}$ and ${}^{-}$. The ${}^{-}$ loss can also be used to target ${}^{\circledR}$ by replacing the minimization over $j=1,\dots,p$ with a minimization over $j\in\mathcal{S}_{r}$, which we denote as ${}^{\circledR}$. The experiment section will show that these losses are not only effective in shrinking the average SEV but often result in attaining the minimum possible SEV value of 1 for most or all queries.
One thing to note is that both All-Opt and Vol-Opt assume that the reference is predicted negative; that is, $\mathcal{L}_{f}(\mathbf{0})=0$. Thus, if the reference is predicted as positive, then optimization for ${}^{+}$ leads to a trivial solution where ${}^{+}$$=0$ for all samples. Similarly ${}^{-}$ would not find a shortest path (or any path for that matter) from the observation to the negative class, since the reference is positive. Thus, to ensure the reference is predicted to be negative, we add a term that penalizes the reference receiving a positive prediction:
@@
where $\theta>0$ is a margin parameter, usually $\theta=0.05$. This term is $(\mathcal{T}-\theta)$ as long as the reference is predicted negative. As soon as it exceeds that amount, it is penalized (increasing linearly in $g(\tilde{\boldsymbol{x}})$).
To put these into an algorithm, we optimize a linear combination of different loss terms,
@@ 5 Experiments
In this section, we present experiments on six real-world datasets to demonstrate the applicability of SEV: (i) UCI Adult Income dataset for predicting income levels (Dua and Graff, 2017), (ii) FICO Home Equity Line of Credit Dataset for assessing credit risk, used for the Explainable Machine Learning Challenge (FICO, 2018), (iii) UCI German Credit dataset for determining creditworthiness (Dua and Graff, 2017), (iv) MIMIC-III dataset for predicting patient outcomes in intensive care units (Johnson et al., 2016a, b), (v) the COMPAS dataset (Jeff Larson and Angwin, 2016; Wang et al., 2022) for predicting recidivism, and (vi) the Diabetes dataset (Strack et al., 2014) for predicting if patients will be re-admitted within two years. Additional details on data and preprocessing are in Appendix A.
We train four reference binary classifiers: (i, ii) Logistic Regression classifiers with $\ell_{1}$ (L1 LR) and $\ell_{2}$ (L2 LR) penalties, (iii) a Gradient Boosting Decision Tree classifier (GBDT), and (iv) a 2-layer Multi-Layer Perceptron (MLP). In addition, we train “optimized" variants of these models, in which the SEV penalties described in the previous sections are implemented. Details on training the methods are in Appendix B. Below, we report the average test SEV; i.e. averaged across queries in the test set.
@@
We first compute ${}^{+}$ and ${}^{-}$ for our four reference classifiers. We emphasize that these models are not optimized for SEV, as they are standard machine-learning models. All models were run for 10 different train-test splits and the mean and standard deviation across these splits of the average test Accuracy, AUC, ${}^{+}$ and ${}^{-}$ on the German Credit data are shown in Table 1. Results for the other datasets are listed in Appendix C.
@@
Even though the German Credit models for L2LR, MLP and GBDT used up to all 20 features (they are not sparse), we observe from Table 1 that most of the queries need only to align one reference feature value with the query value to make positive prediction (${}^{+}$), and we need to align fewer than 2 feature values from the query to the reference for a negative prediction. (${}^{-}$). Thus, even without optimization for SEV at all, most models already have sparse explanations. However, there is still room for improvement, as we now show.
@@
While existing models already have low SEVs on real datasets, our SEV optimization algorithm proposed in the previous section can further decrease SEVs – creating accurate models with extreme decision sparsity. To illustrate, Figure 2 shows the impact on ${}^{+}$ of All-Opt and Vol-Opt on a reference linear classifier trained on the Adult dataset. Figure 1(a) shows the SEV distribution across queries; while L1 LR and L2 LR have reasonably low SEVs (per the discussion in the previous section), both of the proposed optimization algorithms are able to find models for which the ${}^{+}$ of all queries is 1 (the minimum possible SEV). Table 4 in the Appendix shows that this incurs negligible performance loss. Figure 1(b) shows how the optimization affects the ${}^{+}$ of each queries. Its biggest impact on the mean is to reduce the explanations of many queries from involving four terms to a single one, with no query’s explanation becoming more complicated. Additionally, the predicted class distribution is comparable, and the impact of the optimization is not simply to predict a greater number of negative labels; it actually targets the explanation sparsity.
@@
The L1 LR, L2 LR, ${}^{+}$ LR, and Vol-Opt LR methods are all compared in Table 2. L1 LR has already attained low SEVs, possibly because its models are globally sparse. Nevertheless, even without global sparsity, L2 penalized linear regression also generates a low SEV model. This suggests that if our objective is to construct models with low SEVs, we should not just focus on producing globally sparse models, as this may be unnecessarily restrictive. Moreover, after ${}^{+}$ optimization, neither the ${}^{+}$ LR nor the Vol-Opt LR optimization approaches produce sparse models. Thus, it is not necessary to have global sparsity to achieve sparse explanations.
@@ 5.4 Explaining a fixed model: comparing SEV to other local explanation methods
Here, we use GBDT to generate a model, and compare ${}^{-}$ with various post hoc explanation methods, specifically: SHAP-C (using treeSHAP and kernelSHAP) (Lundberg and Lee, 2017; Ramon et al., 2020), LIME-C (Ribeiro et al., 2016; Ramon et al., 2020), and Diverse Counterfactual Explanations (DiCE) (Mothilal et al., 2020). LIME and SHAP are post-hoc explainers that provide local feature importance scores to the prediction. The idea of SHAP-C and LIME-C is to order the features by importance, and based on that ordering, progressively (from the most important feature to less important ones) align the queries’ features with the reference values. We then calculate the number of alignments needed for the point to change from being positively predicted to negatively predicted (shown as Flip Number in Figure 4). We also compare our methods to DiCE, counting the number of features it needs to be modify to generate a counterfactual explanation.
(a) Flip Numbers and Time Consumption
@@
Figure 4 shows that on average LIME-C and SHAP-C flip about half of the features in the model (which uses 7 features in total) and do not induce sparse explanations. This conclusion aligns with Fernández-Loría et al. (2022)’s observation that high local importance features may not always influence model decisions.
@@
Even though DiCE applies backward feature selection after generating explanations generation to control sparsity (Mothilal et al., 2020), Figure 4 shows that ${}^{-}$ provides sparser explanations than DiCE. Moreover, DiCE provide less interpretable explanations than ${}^{-}$. Recall that DiCE uses minimal edit distance to generate counterfactual explanations, whereas we find a minimal number of flips between original values and reference values. As we discussed in Related Works, the edit-distance-based counterfactual explanations may not be as interpretable to humans (Delaney et al., 2023) because the counterfactual often ends up in a low-density region where one cannot determine the class with certainty. The counterfactual explanations generated by DiCE can even lead away from the general direction of the other class. For instance, we found that on a criminal recidivism dataset, DiCE stated that a high-risk criminal offender who commits more crimes would be classified as lower risk, which is paradoxical. The reason for the unreasonable explanation is that DiCE can reference any negatively-predicted point, including outliers. In contrast, the reference point for SEV is the center of the negative population, thus the direction of the SEV is generally meaningful: towards the negatives. For example, the reference value used in our approach for the number of prior misdemeanors is 0, which is more natural. More examples for the comparison between DiCE and ${}^{-}$ are in Appendix E. In addition, Figure 3(a) shows that DiCE takes more time to generate stable sparse counterfactual explanations than ${}^{-}$. The runtime for all datasets can be seen in Appendix D.2, and ${}^{-}$ shows competitive runtimes.
@@ 5.5 Restricted SEV in practice
In this restricted setting, there may exist queries whose predictions do not flip even if all nonrestricted features are flipped. We call these queries “unexplainable.” Figure 4(a) shows the proportion of unexplainable queries in the restricted setting and their relevant ${}^{\circledR}$ before and after restriction and optimization of the ${}^{\circledR}$ objective. Figure 4(b) shows the proportion of queries using each feature for sparse explanations before (Red) and after restriction and optimization (Blue). After introducing restricted features, over one-third of the queries are unable to be explained by the baseline model. However, after ${}^{\circledR}$ is applied, the percentage of unexplainable queries decreases to 3% without much performance loss. Moreover, Figure 4(b) shows that after applying ${}^{\circledR}$ with age and sex=female as restricted features, prior_count becomes the most important factor for explanations for almost all queries, which is more actionable. In real cases, we might want to replace this feature with a dynamic feature, e.g., the number of priors over the last 5 years, to allow individuals to reduce that feature over time through their actions. In that case, one’s explanation would often be simply: you are predicted as high risk because you committed a crime in the last 5 years. More results on the performance of ${}^{\circledR}$ are shown in Appendix F. These results demonstrate that ${}^{\circledR}$ works well in restricted cases.

### 09 Mar 2024 v3

@@ 4.2 Individual-based SEV Optimization
The volume-based SEV loss above approximates queries as arising uniformly over $\mathcal{X}$, which might not be true. Making the less restrictive assumption that the distribution of a training dataset is close to that of test data, we can instead focus on minimizing the average training SEV. The difficulty here lies in the combinatorial number of steps required to calculate SEV. However, we have found that most trained classifiers yield low average SEVs (roughly $<3$) on real datasets. With the majority of training data having low SEV, we can design the optimization objective to penalize queries with ${}^{+}$ not equal to 1. On the other hand, we maximize the fraction of points where ${}^{+}$ is exactly 1. Given a class $\mathcal{F}$ of classification models that output the probability of belonging to the positive class, the loss term for minimizing ${}^{+}$ on a training set is:
@@
$-\frac{1}{n^{+}}\sum_{i=1}^{n^{+}}\min\left(\max_{j=1,\dots,p}g(\boldsymbol{e}_{j}\odot\boldsymbol{x}_{i}+(\mathbf{1}-\boldsymbol{e}_{j})\odot\tilde{\boldsymbol{x}}),~{}\mathcal{T}\right)$
where $\boldsymbol{x}_{i}$ is the $i$-th query in the training data, $n^{+}$ represents the number of positively predicted queries in the dataset, and $\boldsymbol{e}_{j}$ denotes the vector with a 1 in the $j^{th}$ coordinate and 0’s elsewhere, and $g:\mathbb{R}^{p}\to[0,1]$ is the continuous estimator associated with classifier $f$ (e.g., in logistic regression, $g$ is a sigmoid function), and $\mathcal{T}$ is the threshold value for classification, which is usually 0.5 for balanced data. Here, we consider all vertices 1 step away from the reference and see whether at least one of them is predicted as positive (i.e., ${}^{+}$=1). The min operation taken with the threshold $\mathcal{T}$ is used to fix the value of the objective when the prediction is positive since we do not care where on the positive side of the decision boundary a prediction is, as long as it is positive. Similarly, we propose the following loss term for minimizing ${}^{-}$:
@@
We call these individual-based SEV losses ${}^{+}$ and ${}^{-}$. The ${}^{-}$ loss can also be used to target ${}^{\circledR}$ by replacing the minimization over $j=1,\dots,p$ with a minimization over $j\in\mathcal{S}_{r}$, which we denote as ${}^{\circledR}$. The experiments section will show that these losses are not only effective in shrinking the average SEV but often result in attaining the minimum possible SEV value of 1 for most or all queries.
One thing to note is that both All-Opt and Vol-Opt assume that the reference is predicted negative; that is, $\mathcal{L}_{f,\textrm{query}}(\mathbf{0})=0$. Thus, if the reference is predicted as positive, then optimization for ${}^{+}$ leads to a trivial solution where ${}^{+}$$=0$ for all samples. Similarly ${}^{-}$ would not find a shortest path (or any path for that matter) from the observation to the negative class, since the reference is positive. Thus, to ensure the reference is predicted to be negative, we add a term that penalizes the reference receiving a positive prediction:
@@
where $\theta>0$ is a margin parameter, for example, $\theta=0.05$. This term is $(\mathcal{T}-\theta)$ as long as the reference is predicted negative. As soon as it exceeds that amount, it is penalized (increasing linearly in $g(\tilde{\boldsymbol{x}})$).
Combining the loss terms above, we optimize a linear combination of them,
@@ 5 Experiments
We present experiments on six real-world datasets: (i) the UCI Adult Income dataset for predicting income levels (Becker and Kohavi, 1996; Dua and Graff, 2017), (ii) the FICO Home Equity Line of Credit Dataset for assessing credit risk, used for the Explainable Machine Learning Challenge (FICO, 2018), (iii) the UCI German Credit dataset for determining creditworthiness (Hofmann, 1994; Dua and Graff, 2017), (iv) the MIMIC-III dataset for predicting patient outcomes in intensive care units (Johnson et al., 2016a, b), (v) the COMPAS dataset (Jeff Larson and Angwin, 2016; Wang et al., 2022) for predicting recidivism, and (vi) the Diabetes dataset (Strack et al., 2014; Dua and Graff, 2017) for predicting if patients will be re-admitted within two years. Additional details on data and preprocessing are in Appendix A.
We train four reference binary classifiers: (i, ii) Logistic Regression classifiers with $\ell_{1}$ (L1 LR) and $\ell_{2}$ (L2 LR) penalties, (iii) a Gradient Boosting Decision Tree classifier (GBDT), and (iv) a 2-layer Multi-Layer Perceptron (MLP). In addition, we train “optimized” variants of these models, in which the SEV penalties described in the previous sections are implemented. Details on training the methods are in Appendix B. Below, we report the average test SEV; i.e. averaged across queries in the test set.
@@
We first compute ${}^{+}$ and ${}^{-}$ for our four reference classifiers. We emphasize that these models are not optimized for SEV, they were created with standard machine-learning algorithms. All algorithms were run for 10 different train-test splits and the mean and standard deviation across these splits of the average test accuracy, AUC, ${}^{+}$ and ${}^{-}$ on the German Credit data are shown in Table 1. Results for the other datasets are listed in Appendix C.
@@
Even though the German Credit models for L2LR, MLP and GBDT used up to all 20 features (they are not sparse), we observe from Table 1 that most of the queries need only to align one reference feature value with the query value to change to a positive prediction (${}^{+}$), and we need to align fewer than 2 feature values from the query to the reference for a negative prediction (${}^{-}$). Thus, even without optimization for SEV at all, most models already have sparse explanations. However, there is still room for improvement, as we now show.
@@
While existing models already have low SEVs on real datasets, our SEV optimization algorithm proposed in the previous section can further decrease SEVs – creating accurate models with extreme decision sparsity. To illustrate, Figure 2 shows the impact on ${}^{+}$ of All-Opt and Vol-Opt on a reference linear classifier trained on the Adult dataset. Figure 1(a) shows the SEV distribution across queries; while L1 LR and L2 LR have reasonably low SEVs (per the discussion in the previous subsection), both of the proposed optimization algorithms are able to find models for which the ${}^{+}$ of all queries is 1 (the minimum possible SEV). Table 7 in the Appendix shows that this incurs negligible performance loss. Figure 1(b) shows how the optimization affects the ${}^{+}$ of each query. Its biggest impact on the mean is from reducing the explanations of many queries from involving 2 to 5 terms to a single one, with no query’s explanation becoming more complicated. Additionally, the predicted class distribution is comparable, and the impact of the optimization is not simply to predict a greater number of negative labels; it actually targets the explanation sparsity.
@@
Moreover, to evaluate the time consumption of our methods, we have tried different dataset sizes. The detailed information for each dataset is shown in Appendix A. To better compare the training time consumption, we have used MLP as the baseline model and compared its running time with and without All-Opt for different datasets. All the models were run 10 times and we report the mean and standard deviation in Table 2.
Dataset Baseline ${}^{+}$ ${}^{-}$
COMPAS $10.5\pm 0.69$ $26.9\pm 1.06$ $21.7\pm 0.03$
Adult $61.1\pm 0.15$ $164\pm 1.56$ $136\pm 1.22$
MIMIC III $78.2\pm 1.09$ $207\pm 2.71$ $167\pm 2.42$
German Credit $1.81\pm 0.01$ $5.01\pm 0.01$ $4.08\pm 0.04$
FICO $16.9\pm 0.78$ $47.6\pm 0.34$ $37.4\pm 1.30$
Diabetes $188\pm 3.08$ $606\pm 3.65$ $517\pm 3.31$
Table 2: The training time (seconds) comparison with and without All-Opt in MLP
@@
The L1 LR, L2 LR, ${}^{+}$ LR, and Vol-Opt LR methods are all compared in Table 3. L1 LR has already attained low SEVs, possibly because its models are globally sparse. Nevertheless, even without global sparsity, L2 penalized linear regression also generates a low SEV model. This suggests that if our objective is to construct models with low SEVs, we should not just focus on producing globally sparse models, as this may be unnecessarily restrictive. Moreover, after ${}^{+}$ optimization, neither the ${}^{+}$ LR nor the Vol-Opt LR optimization approaches produce sparse models and yet have perfect SEV’s. Thus, it is not necessary to have global sparsity to achieve sparse explanations.
@@ 5.4 Explaining a fixed model: comparing SEV to other local explanation methods
Here, we use GBDT to generate a model on the COMPAS dataset, and compare ${}^{-}$ with various post hoc explanation methods, specifically: SHAP-C (using treeSHAP and kernelSHAP) (Lundberg and Lee, 2017; Ramon et al., 2020), LIME-C (Ribeiro et al., 2016; Ramon et al., 2020), and Diverse Counterfactual Explanations (DiCE) (Mothilal et al., 2020). LIME and SHAP are post-hoc explainers that provide local feature importance scores to the prediction. The idea of SHAP-C and LIME-C is to order the features by importance, and based on that ordering, progressively (from the most important feature to less important ones) align the queries’ features with the reference values. We then calculate the number of alignments needed for the point to change from being positively predicted to negatively predicted (shown as Flip Number in Figure 4). We also compare our methods to DiCE, counting the number of features it modifies to generate a counterfactual explanation.
(a) Flip Numbers and Mean Explanation Time Consumption
@@
Figure 4 shows that, on average, LIME-C and SHAP-C flip about half of the features in the model (which uses 7 features in total) and do not induce sparse explanations. This conclusion aligns with Fernández-Loría et al. (2022)’s observation that high locally important features may not always influence model decisions.
@@
Even though DiCE applies backward feature selection after generating explanations to control sparsity (Mothilal et al., 2020), Figure 4 shows that ${}^{-}$ provides sparser explanations than DiCE. Moreover, DiCE provides less interpretable explanations than ${}^{-}$. Recall that DiCE uses minimal edit distance to generate counterfactual explanations, whereas we find a minimal number of flips between original values and reference values. As we discussed in Related Works, the edit-distance-based counterfactual explanations may not be as interpretable to humans (Delaney et al., 2023) because the counterfactual often ends up in a low-density region where one cannot determine the class with certainty. The counterfactual explanations generated by DiCE can even lead away from the general direction of the other class. For instance, we found that on a criminal recidivism dataset, DiCE stated that a high-risk criminal offender who commits more crimes would be classified as lower risk, which is paradoxical. This unreasonable explanation arises because DiCE can reference any negatively-predicted point, including outliers. In contrast, the reference point for SEV is the center of the negative population, thus the direction of the SEV is generally meaningful: towards the negatives. For example, the reference value used in our approach for the number of prior misdemeanors is 0, which is more natural.
Table 4 shows counterfactual explanations for two different queries. For the first, DiCE suggests that the number of juvenile crimes should go from 1 to 5 to change the prediction from ‘will recidivate’ to ‘won’t recidivate.’ This is nonsensical, as the number of crimes should be positively correlated with the propensity to recidivate. In contrast, ${}^{-}$ suggests that the number of prior crimes should go from 2 to 1, which makes sense. Similarly, for the second query, DiCE suggests that the number of juvenile crimes should go from 3 to 7 and the number of juvenile misdemeanors from 2 to 11 in order to change the prediction from ‘will recidivate’ to ‘won’t recidivate.’ Again, this explanation is nonsensical. In contrast, ${}^{-}$ provides two possible explanations: (i) that the number of prior crimes should go from 3 to 1 or (ii) that the number of juvenile misdemeanors should go from 2 to 0. Again, this makes more sense. Given that DiCE uses a random initialization to find a counterfactual explanation whereas ${}^{-}$ provides the minimum number of feature alignments to cross the decision boundary, we would expect ${}^{-}$ to be more stable and to generate useful explanations.
Explanations Sex Age
Juv fel
counts
Juv misd
count
Juvenile
crimes
Priors
counts
Current Charge
Degree
Predicted
Label
Query 1 Male 23 0 1 1 2 Not Felony 1
DiCE 5 0
${}^{-}$ 1 0
Query 2 Male 19 1 2 3 3 Felony 1
DiCE 11 7 0
${}^{-}$ 0 0
1 0
Table 4: DiCE and ${}^{-}$ explanations for two different queries in COMPAS
In addition, SEV has shown competitive speed in generating local explanations. Figure 3(a) shows that DiCE takes more time to generate stable sparse counterfactual explanations than ${}^{-}$. Table 5 shows the detailed time required to generate a local explanation for our method and baselines SHAP, LIME, and DiCE, for several datasets. We use GBDT as a baseline model and evaluate the median explanation runtime on queries in the test set. The table shows that ${}^{-}$ has a reasonable running time for black-box models on most datasets. We can observe that the explanation time is better than DiCE in all the datasets since it only flips from feature values to the reference. The reason why the FICO Dataset takes a long time is that its average ${}^{-}$ is about 4, which means that the search in the SEV hypercube starting from the query is too deep. However, if the time consumption for generating explanations is too long, it means that this model cannot be sparsely explained, but we can use All-Opt or Vol-Opt methods to optimize our model. Importantly, other types of explanations (SHAP, LIME, DiCE) do not have optimization methods such as All-Opt or Vol-Opt. The detailed comparison between KernelSHAP and ${}^{-}$ is shown in Appendix D.
Dataset TreeSHAP KernelSHAP LIME DiCE ${}^{-}$
MIMIC $1.13\pm 0.01$ $17.30\pm 0.10$ $6.10\pm 0.05$ Crashed $0.45\pm 0.03$
German
Credit
$0.53\pm 0.00$ $9.70\pm 0.17$ $7.11\pm 0.07$ $26.70\pm 0.91$ $1.37\pm 0.05$
FICO $0.56\pm 0.02$ $8.25\pm 0.27$ $6.05\pm 0.14$ $27.82\pm 0.05$ $15.75\pm 1.01$
Table 5: The median explanation runtime (in $10^{-2}$ s) for query units predicted as positive. The values in the table differ from those in Figure 3(a) since Figure 3(a) reports the mean runtime instead of median.
@@ 5.5 Restricted SEV in practice
In this restricted setting, there may exist queries whose predictions do not flip even if all nonrestricted features are flipped. We call these queries “unexplainable.” Figure 4(a) shows the proportion of unexplainable queries in the restricted setting and their relevant ${}^{\circledR}$ before and after restriction and optimization of the ${}^{\circledR}$ objective. Figure 4(b) shows the proportion of queries using each feature for sparse explanations before (Red) and after restriction and optimization (Blue). After introducing restricted features, over one-third of the queries are unable to be explained by the baseline model. However, after ${}^{\circledR}$ is applied, the percentage of unexplainable queries decreases to 3% without much performance loss. Moreover, Figure 4(b) shows that after applying ${}^{\circledR}$ with age and sex=female as restricted features, prior_count becomes the most important factor for explanations for almost all queries, which is more actionable. In real cases, we might want to replace this feature with a dynamic feature, e.g., the number of priors over the last 5 years, to allow individuals to reduce that feature over time through their actions. In that case, one’s explanation would often be simply: you are predicted as high risk because you committed a crime in the last 5 years. More results on the performance of ${}^{\circledR}$ are shown in Appendix E. These results demonstrate that ${}^{\circledR}$ works well in restricted cases.